OpenStudy (purexyz):

Find the length of a if the measure of angle A equals 62 degrees m the measure of angle B equals 39 degrees and b = 15.

OpenStudy (purexyz):

this is using sine please help

hartnn (hartnn):

then use the sine formula, \(\Large \dfrac{a}{\sin A}= \dfrac{b}{\sin B}\) where A = 62, b =15, B = 39 plug in :)